OpenStudy (anonymous):

Here are Janet's steps for solving the system below by eliminating x. 2x + 3y = -1 5x + 8y = 0 Step One: Multiply the first equation by 5. 10x + 15y = -5 Step Two: Multiply the second equation by -2. -10x - 16y = 0 Step Three: Add the two equations. y = -5 Step Four: Substitute y back into the first equation. 2x + 3(-5) = -1 Step Five: Solve for x. 2x - 15 = -1 2x = 14 x = 7 Is she correct? Justify your answer. If incorrect, describe the error.

OpenStudy (cwrw238):

have a look at step 3 - is the result of adding the 2 equations correct?

OpenStudy (cwrw238):

what is +15y + (-16y) ?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

well its -y so it should be -y = -5 which makes y = 5 right?

OpenStudy (cwrw238):

to you follow that?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

ok so although the rest of the math is correct the value of x cannot be 7 because y is not = -5 can you find the correct value of x ?

OpenStudy (cwrw238):

i'll start you ogg 2x + 3(5) = -1

OpenStudy (cwrw238):

*off lol!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

haha i figured so

OpenStudy (anonymous):

honestly im not very experienced on this

OpenStudy (cwrw238):

2x + 15 = -1 the aim is to get the terms in x on one side of the '=' and numbers on the right first subtract 15 from both sides: 2x + 15 - 15 = -1 - 15 (15 - 15 is of course = 0) so 2x = -16 now divide both sides by 2 to find x 2x /2 = -16/2 x = -8

OpenStudy (anonymous):

and y=5

OpenStudy (anonymous):

so shes incorrect

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

sweet, thank you for the help :)

OpenStudy (cwrw238):

there is only one mistake - step 3 but this makes both values wrong

OpenStudy (cwrw238):

yw